A 550 mL sample of nitrogen gas is warmed from 77C to 86C. Find its new volume if the pressure remains constant.



1
Expert's answer
2021-12-20T05:24:02-0500

V1= 550ml

V2 =?

T1= 77°C +273 = 350K

T2 = 86°C +273 =359K


"\frac{V1}{T1}= \frac{V2}{T2}"


"=\frac{550ml}{350K} = \frac{V2}{359K}"


"V2 =\frac{550ml ×359K}{350K} =564.14 ml"